OpenStudy (anonymous):

Can someone explain to me why pi/3 + 2pi is 7pi/3? How in the world does that work?

ganeshie8 (ganeshie8):

think of pi as pie. how many pi/3's are there in 2pi ?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Oh my gosh, I'm an idiot. You just find a common denominator. Wow. Thanks guys. I'm so slow. God bless you all! :)

OpenStudy (anonymous):

for the same reason that \[2+\frac{1}{3}=\frac{7}{3}\]

OpenStudy (anonymous):

latex not working?

OpenStudy (ybarrap):

Nope, not an idiot. Human -- we are all. :)

OpenStudy (ybarrap):

Latex is not working neither is Draw.

OpenStudy (anonymous):

Haha. Thanks. I just feel really dumb sometimes. :P Yes, I noticed that. :c

OpenStudy (anonymous):

How do I do arcsin?

OpenStudy (anonymous):

Because we use pi as in a pro numeral. Treat it like all other variables. Imagine pi as it's...the letter x. So, x/3 + 2x Now you just add the fraction like you would, find the least common denominator, or 3. x/3 + 6x/3 = 7/3x You can just replace x with pi. Same thing. Lol. x

OpenStudy (anonymous):

Thank you. :) Do you know how to do arcsin(sqrt3/2)?